1. Determine the poles and residue at each poles of

f(z)=2z+1/z²-z-2 over C=|z|=5/2 Hence Evaluate

∮c 2z+1/z²-z-2 dz over C=|z|=5/2


1
Expert's answer
2021-06-08T13:20:17-0400

"z\u00b2-z-2=(z-2)(z+1)"


Poles:

"z_0=2,z_0=-1"


"res\ f(2)=\displaystyle{\lim_{z\to 2}}(f(z)(z-2))=\displaystyle{\lim_{z\to 2}}(\frac{2z+1}{z+1})=5/3"


"res\ f(-1)=\displaystyle{\lim_{z\to -1}}(f(z)(z+1))=\displaystyle{\lim_{z\to -1}}(\frac{2z+1}{z-2})=1/3"


"f(z_0)=\frac{1}{2\pi i}\oint_C\frac{f(z)}{z-z_0}dz" , "z_0\isin D"


"\frac{1}{(z-2)(z+1)}=\frac{A}{z-2}+\frac{B}{z+1}"


"A(z+1)+B(z-2)=1"

"A+B=0"

"A-2B=1"

"B=-1\/3,A=1\/3"


"\oint_C\frac{f(z)}{z-z_0}dz=\frac{2\pi i}{3}\oint\frac{2z+1}{z-2}dz-\frac{2\pi i}{3}\oint\frac{2z+1}{z+1}dz=\frac{2\pi i}{3}(2z+1)|_{z=2}-\frac{2\pi i}{3}(2z+1)|_{z=-1}="


"=\frac{10\pi i}{3}+\frac{2\pi i}{3}=4\pi i"